# One-Layer Transformers are Provably Optimal for In-context Reasoning and Distributional Association Learning in Next-Token Prediction Tasks

### v2

@@ 3.2 Training Dynamics and On-Convergence Behavior: A Distributionally Agnostic Analysis
i.e. the gradient vectors are normalized by their Euclidean norm. A similar normalized gradient descent update for transformers has been used in Huang et al. (2024b) for learning an injective map on the vocabulary $[N]$. The following theorem shows that the update rule (5) can be implemented without the knowledge of the distribution of $z$. Moreover, the population loss converges at a linear rate to zero. The proof can be found in Appendix A.2.
@@
Starting from $\lambda_{q,0}=0$ for all $q\in{\mathcal{Q}}$, the update rule (5) is equivalent to $|\lambda_{q,t}=\frac{\eta t}{\absolutevalue{{\mathcal{Q}}}}$ for all $t\geq 1$. Moreover, $L(\bm{\lambda}_{t})=O(N\exp(-t))$.
Generalization to unseen output words. The only correct strategy for solving the noiseless data model in Definition 2.1 is to predict the word that comes after a trigger token. That is, the position of the trigger token is the only important factor in this task. Such a strategy would easily generalize to a sentence which contains a new bigram $(q,y_{\mathrm{test}})$, where $y_{\mathrm{test}}$ is a non-trigger non-output word. We characterize this generalization ability of our parameterized transformers in the following theorem.
@@
In particular, this implies that $\Pr[z_{H+1}=y_{\mathrm{test}}\mid\bm{\lambda}_{t}]\geq 0.5$ for $t\geq\ln(N)/\eta$.
Reparameterization versus Directional Convergence. In addition to the convergence of the loss function, existing works (e.g. Ji and Telgarsky, 2021; Huang et al., 2024b) on the training dynamics of neural networks learned with cross-entropy loss have shown that the trainable matrices directionally convergence to an optimal solution. More formally, a sequence of ${\bm{A}}_{t}$ directionally converges to some $\bm{A}$$_{*}$ if $\lim_{t\to\infty}\langle\frac{{\bm{A}}_{t}}{\norm{{\bm{A}}_{t}}},\frac{{\bm{A}}_{*}}{\norm{{\bm{A}}_{*}}}\rangle=1$.
In our work, the joint query-key matrix $\bm{W}$} is reparameterized as a form of associative memory of the trigger tokens, rather than emerging from runing gradient descent for minimizing the population loss $L$, as in  (Bietti et al., 2023; Chen et al., 2025) This raises a natural question: if we use the reparameterization on $\bm{U}$} and $\bm{V}$} but not $\bm{W}$}, will running gradient descent on $\bm{W}$} directionally converges to $\sum_{q}E(q)\tilde{E}^{\top}(q)$? Theorem 3.4 (full proof in Appendix A.4) provides a negative answer.
@@
This result shows that the space of optimal solutions for this in-context reasoning problem is much more complicated than that of the max-margin problems trained with logistic loss in Ji and Telgarsky (2021); Huang et al. (2024b). Moreover, directional convergence cannot fully characterize the set of optimal solutions.
4 Noisy Learning ($\alpha>0$)
Recall that our noisy data model is a variant of the setting in Chen et al. (2025). Similar to Chen et al. (2025), we assume that there is only one trigger token $q$, i.e. ${\mathcal{Q}}=\{q\}$. Unlike Chen et al. (2025), we allow any distributions of bigrams $(q,y)$ and $(q,\tau)$ in the sentence, and do not assume that the ratio of the frequencies of the bigrams $(q,\tau)$ and $(q,y)$ is $\frac{\alpha}{1-\alpha}$. In other words, our results are based on the distribution of the label $z_{H+1}$ instead of the distribution of the bigrams $(q,\tau)$ and $(q,y)$. We emphasize that a Bayes-optimal solution for our noisy data model will also be a Bayes-optimal solution for the model in Chen et al. (2025).
First, we show that by relying on just the distribution of $z_{H+1}$, our one-layer transformer in (1) is capable of making the population loss arbitrarily close to the loss of the Bayes optimal strategy. We start by characterizing the Bayes optimal strategy. Fix an output word $y$. The label of a sentence that contains both $(q,\tau)$ and $(q,y)$ is either $\tau$ with probability $\alpha$ or $y$ with probability $1-\alpha$, independent of other tokens in the sentence. Thus, the Bayes optimal strategy is to predict $\tau$ and $y$ with the same probabilities. Let $\hat{y}$ be the prediction of this strategy. Its expected loss is equal to the entropy
$L_{\mathrm{Bayes}}$ $=\Pr(z_{H+1}=\tau)(-\ln(\Pr_{\mathrm{Bayes}}(\hat{y}=\tau))+\sum_{y\in{\mathcal{O}}}\Pr(z_{H+1}=y)(-\ln(\Pr_{\mathrm{Bayes}}(\hat{y}=y))$
$=-\alpha\ln\alpha-(1-\alpha)\ln(1-\alpha).$
The following lemma shows that using linear and ReLU attentions, applying softmax on $\xi=\xi_{A}+\xi_{F}$ leads to a Bayes optimal strategy as $\lambda\to\infty$.
Lemma 4.1.
Let $y\in{\mathcal{O}}$ be arbitrary and $\lambda,\gamma\in\mathbb{R}$. Under Assumption 2.2, by setting ${\bm{U}}=[E(1)~E(2)~\dots~E(N)~E(N+1)]^{\top},{\bm{V}}={\bm{I}}_{d},{\bm{W}}=\lambda(E(q)\tilde{E}^{\top}(q)-E(q)E^{\top}(\tau))$ and ${\bm{F}}=E(\tau)(\gamma E^{\top}(q)+\tilde{E}^{\top}(q))$, for all $\alpha\in(0,1)$, for both linear and ReLU attention, we obtain
Remark 4.2.
Lemma 4.1 is distribution-agnostic in the sense that it holds for any word distribution as long as the three conditions of the data model are satisfied. Moreover, our result holds even for $\alpha>0.5$. This is a major advantage over the existing results in Chen et al. (2025), which required $\alpha\leq 0.5$. Setting $\alpha>0.5$ also reflects a wider range of practical scenarios where the generic bigrams such as “of the” often appear more frequently than context-dependent bigrams such as “of Bob”.
@@ 4.2 Training Dynamics and On-Convergence Behavior: A Finite-Sample Analysis
Similar to Section 3.2, we focus on the analysis of linear attention. Let $M$ denote the size of a dataset of $M$ i.i.d sentences $z_{H+1}$ generated from the data model in Definition 2.1. Instead of minimizing the full population loss as in existing works (Bietti et al., 2023; Huang et al., 2024a; Chen et al., 2025), which would require either knowing $\alpha$ or taking $M\to\infty$, we aim to derive a finite-sample analysis that holds for finite $M$ and unknown $\alpha$.
Before presenting our training algorithm and its finite-sample analysis, we first discuss the easier case where $\alpha$ is known and explain why it is difficult to derive the convergence rate of the population loss in (25). Observe that the function $f_{C}(\lambda,\gamma)=-C\lambda-\alpha\gamma+\ln(e^{C\lambda}+e^{C\lambda+\gamma}+N-1)$ is jointly convex and 1/2-smooth with respect to $\lambda$ and $\gamma$. Hence, at first glance, it seems that the convergence rate of $L(\lambda,\gamma)=\mathbb{E}_{C}[f_{C}(\lambda,\gamma)]$ follows from existing results in (stochastic) convex and smooth optimization (Nemirovski et al., 2009). However, as a convex function on unbounded domain with negative partial derivatives, no finite minimizer exists for $L(\lambda,\gamma)$. This implies that minimizing $L(\lambda,\gamma)$ is a multi-dimensional convex optimization problem on astral space (Dud´ık et al., 2022). To our knowledge, nothing is known about the convergence rate to the infimum for this problem.
@@
$L_{\mathrm{emp}}(\lambda)$ $=\frac{1}{M}\sum_{m=1}^{M}-\ln\frac{\exp(\xi_{z_{H+1}}^{(m)})}{\sum_{j\in[N+1]}\exp(\xi_{j}^{(m)})}.$ (7)
Then, we run normalized gradient descent on $L_{\mathrm{emp}}(\lambda)$ with a constant learning rate $\eta>0$. The formal procedure is given in Algorithm 1 in Appendix B.2. The convergence of the population loss of this algorithm is stated in the following theorem.
@@
With probability at least $1-\delta$ over the training set of size $M$, after $t$ iterations of normalized gradient descent on (7), Algorithm 1 guarantees that
$L(\lambda_{t},\hat{\gamma})\leq L_{\mathrm{Bayes}}+\frac{1}{\min(\alpha,1-\alpha)-\sqrt{\frac{\ln(2/\delta)}{2M}}}\frac{\ln(2/\delta)}{2M}+(N-1)e^{-\eta t}.$
Generalization to unseen output words. Theorem 4.3 indicates that the population loss converges to the Bayes-risk as $M$ and $t$ increases. Next, the following theorem shows that Algorithm 1 produces a trained model that generalizes to an unseen output word $\hat{y}\notin{\mathcal{O}},\hat{y}\neq q$. The proof is in Appendix B.3.
@@
$\Pr[z_{H+1}=\hat{y}\mid\lambda_{t},\hat{\gamma}$] $\vcentcolon=\frac{\exp(\xi_{y_{\mathrm{test}}})}{\sum_{j\in[N+1]}\exp(\xi_{j})}=1-\alpha+O\left(\sqrt{\frac{\ln(1/\delta)}{M}}+N^{2}e^{-2\eta t}\right),$
@@
The following theorem shows that after a sufficiently large number of iterations, Algorithm 1 returns a model that satisfies condition (8) for any input sentence generated from our data model.
@@
Algorithm 1 guarantees that with probability at least $1-\delta$, after $|t\geq\max(1,\frac{1}{\eta}\absolutevalue{\ln(1-\alpha+\sqrt{\frac{\ln(2/\delta)}{2M}})-\ln(\alpha-\sqrt{\frac{\ln(2/\delta)}{2M}})}$ iterations, the condition (8) holds for any sentence generated from the noisy data model.

### v3

@@ 3.2 Convergence rate, Generalization and Implicit Bias of Gradient Descent
i.e. the gradient vectors are normalized by their Euclidean norm. A similar NGD update was used in Huang et al. (2024b) for learning an injective map on the vocabulary. The following theorem shows that the update  (3) can be implemented without the knowledge of the distribution of $z$. Moreover, the population loss converges at a linear rate to zero. The proof can be found in Appendix B.2.
@@
Starting from $\lambda_{q,0}=0$ for all $q\in{\mathcal{Q}}$, the update rule (3) is equivalent to $|\lambda_{q,t}=\frac{\eta t}{\absolutevalue{{\mathcal{Q}}}}$ for all $t\geq 1$. Moreover, $L(\bm{\lambda}_{t})=O(N\exp(-\eta t))$.
OOD Generalization to unseen output words. The human-like strategy for solving the noiseless data model in Definition 2.1 is to predict the word that comes after a trigger token. That is, the position of the trigger token is the only important factor in this task. Such a strategy does not depend on what the actual output word $y$ is, and hence would easily generalize to a out-of-distribution sentence where the bigram $(q,y)$ is replaced by a new bigram $(q,y_{\mathrm{test}})$, where $y_{\mathrm{test}}$ is a non-trigger non-output word. The following theorem formalizes this intuition, indicating that our parameterization is precisely implementing this human-like strategy.
@@
In particular, this implies that $\lim_{t\to\infty}\Pr[z_{H+1}=y_{\mathrm{test}}\mid\bm{\lambda}_{t}]=1$.
Reparameterization versus Directional Convergence. In addition to the convergence of the loss function, existing works (e.g. Ji and Telgarsky, 2021; Huang et al., 2024b) on the training dynamics of neural networks learned with cross-entropy loss have shown that the trainable matrices directionally convergence to an optimal solution. More formally, a sequence of ${\bm{A}}_{t}$ directionally converges to some $\bm{A}$$_{*}$ if $\lim_{t\to\infty}\langle\frac{{\bm{A}}_{t}}{\norm{{\bm{A}}_{t}}},\frac{{\bm{A}}_{*}}{\norm{{\bm{A}}_{*}}}\rangle=1$. In our work, the joint query-key matrix $\bm{W}$} is reparameterized as a form of associative memory of the trigger tokens, rather than emerging from runing gradient descent for minimizing the population loss $L$, as in  (Bietti et al., 2023; Chen et al., 2025). This raises a natural question: if we use the reparameterization on $\bm{U}$} and $\bm{V}$} but not $\bm{W}$}, what is the implicit bias of running gradient descent on $\bm{W}$}? In Theorem 3.4 (full proof in Appendix B.4), we show that gradient desecent does not directionally converges to $\sum_{q}E(q)\tilde{E}^{\top}(q)$.
@@
4 Noiseless Learning with Softmax Attention
Fix a sentence with trigger word $q$ and output word $y$. Recall that in order to achieve a zero population logistic loss, it is necessary that the logits $\xi_{y}$ of the output word $y$ approaches infinity. The proof sketch of Lemma 3.1 the previous section showed that having an unbounded attention scores $\sigma({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})$ at $z_{h-1}=q$ naturally allows $\xi_{y}$ to approach infinity. While this unboundedness hold for linear and ReLU attention, it does not hold for softmax attention, where the attention scores are bounded in the range $[0,1]$. The following lemma (proof in Appendix C) shows a modified parameterization that can drive the logits $\xi_{y}$ to infinity under softmax attention.
Lemma 4.1.
Let $s\in\mathbb{R}$ and $\bm{\lambda}=\{\lambda_{k}\in\mathbb{R}_{+}:k\in{\mathcal{Q}}\}$ be a set of $|\absolutevalue{{\mathcal{Q}}}$ of non-negative values. By setting ${\bm{U}}=[E(1)~E(2)~\dots~E(N)]^{\top},{\bm{V}}=s{\bm{I}}_{d}$ and ${\bm{W}}=\sum_{k\in{\mathcal{Q}}}\lambda_{k}E(k)\left(\tilde{E}(k)^{\top}-\sum_{x=1,x\neq k}^{N}\tilde{E}(x)^{\top}\right)$, for softmax attention, we obtain $\lim_{s\to\infty}\lim_{(\lambda_{k})_{k\in{\mathcal{Q}}}\to\infty}L\left({\bm{V}},{\bm{W}}\right)=0.$
This new parameterization has two modifications compared to the one in Lemma 3.1: a subtraction $\sum_{x=1,x\neq k}^{N}\tilde{E}(x)^{\top}$ from $\bm{W}$}, and a new scaling factor $s\in\mathbb{R}$ in the value matrix $\bm{V}$}. The first modification ensures that ${\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h}$ tends to positive and negative infinity for positions $h$ where $z_{h-1}=q$ and $z_{h-1}\neq q$, respectively. Correspondingly, the attention scores $\sigma({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})$ tends to $1$ and 0 for these two cases. The second modification effectively shifts the scaling in $\xi$ from $\bm{W}$} to $\bm{V}$}, and implies that the desired optimality comes from $\lim_{s\to\infty}s\cdot 1=\infty$ and $\lim_{s\to\infty}s\cdot 0=0$.
Note that in the statement of Lemma 4.1, the order of the two limit operations are strict and not exchangable. Therefore, it is an important question that whether this optimality can actually be realized by running normalized gradient descent on $L({\bm{V}},{\bm{W}})$. The following result answers this question in the positive, showing that from a certain initialization, both $s$ and $(\lambda_{k})_{k}$ approaches infinity. Moreover, the population loss $L({\bm{V}},{\bm{W}})$ converges to 0 at a linear rate.
Theorem 4.2.
@@ 4 Noiseless Learning with Softmax Attention
The proof of Theorem 4.2 divides the training process into two distinct phases: $t\leq T_{0}$ and $t>T_{0}$. During this first phase, the signs of the derivatives $\frac{\partial L}{\partial s_{t}}$ may oscillate between +1+1 and 1-1 while $\lambda_{q,t}$ and the attention scores grow quickly. In the second phase, $\lambda_{q,t}$ has become sufficiently large so that the attention scores become more stable, allowing $s_{t}$ to increase monotonically. Similar stage-wise convergence analysis of transformers with softmax attention has been observed before in other tasks such as regression (Huang et al., 2024a) and binary classification (Huang et al., 2025).
5 Noisy Learning
[…]
First, we show that by relying on just the distribution of $z_{H+1}$, the model defined in (1) is capable of making the population loss arbitrarily close to the loss of the Bayes optimal strategy. Fix a trigger word $q$ and an output word $y$. The label of a sentence that contains both $(q,\tau)$ and $(q,y)$ is either $\tau$ with probability $\alpha$ or $y$ with probability $1-\alpha$, independent of other tokens in the sentence. Thus, the Bayes optimal strategy is to predict $\tau$ and $y$ with the same probabilities. Let $\hat{y}$ be the prediction of this strategy. Its expected loss is equal to the entropy $L_{\mathrm{Bayes}}=-\alpha\ln\alpha-(1-\alpha)\ln(1-\alpha).$ Similar to the previous sections, we set and fix the unembedding layer ${\bm{U}}=[E(1)~E(2)~\dots~E(N)~E(N+1)]^{\top}$. The following two lemmas demonstrate that the Bayes optimality of specific parameterization for linear/ReLU and softmax attentions.
Lemma 5.1.
Let $\lambda,\gamma\in\mathbb{R}$. By setting ${\bm{V}}={\bm{I}}_{d},{\bm{W}}=\lambda\sum_{q\in{\mathcal{Q}}}E(q)(\tilde{E}^{\top}(q)-E^{\top}(\tau))$ and ${\bm{F}}=E(\tau)(\sum_{q\in{\mathcal{Q}}}\gamma E^{\top}(q)+\tilde{E}^{\top}(q))$ and using linear or ReLU attention, we obtain
$\lim_{\lambda\to\infty,\gamma\to\ln\frac{\alpha}{1-\alpha}}L(\lambda,\gamma)\vcentcolon=\lim_{\lambda\to\infty,\gamma\to\ln\frac{\alpha}{1-\alpha}}\mathbb{E}_{y,z_{1:H+1}}\left[-\ln\frac{\exp(\xi_{z_{H+1}})}{\sum_{j\in[N+1]}\exp(\xi_{j})}\right]=L_{\mathrm{Bayes}}.$
Lemma 5.2.
Let $s,\lambda,\gamma\in\mathbb{R}$. By setting ${\bm{V}}=s{\bm{I}}_{d},{\bm{W}}=\lambda\sum_{q\in{\mathcal{Q}}}E(q)(\tilde{E}^{\top}(q)-2E^{\top}(\tau)-\sum_{x=1,x\neq q}^{N}\tilde{E}(x)^{\top})$ and ${\bm{F}}=E(\tau)(\sum_{q\in{\mathcal{Q}}}\gamma E^{\top}(q)+\tilde{E}^{\top}(q))$ and using softmax attention,
$\lim_{s\to\infty,\lambda\to\infty,\gamma\to\ln\frac{\alpha}{1-\alpha}}L(s,\lambda,\gamma)\vcentcolon=\lim_{s\to\infty}\lim_{\lambda\to\infty,\gamma\to\ln\frac{\alpha}{1-\alpha}}\mathbb{E}_{y,z_{1:H+1}}\left[-\ln\frac{\exp(\xi_{z_{H+1}})}{\sum_{j\in[N+1]}\exp(\xi_{j})}\right]=L_{\mathrm{Bayes}}.$
@@ 5.1 Approximation Capabilities of One-Layer Transformers On Noisy Setting
These results are distribution-agnostic in the sense that it holds for any word distribution as long as the conditions of the data model are satisfied. Moreover, they hold even for $\alpha>0.5$. This is a major advantage over the existing results in Chen et al. (2025), which required $\alpha\leq 0.5$. Setting $\alpha>0.5$ also reflects a wider range of practical scenarios where the generic bigrams such as “to the” often appear more frequently than context-dependent bigrams such as “to Bob”.
@@ 5.2 Training Dynamics and On-Convergence Behavior: A Finite-Sample Analysis
For ease of exposition, we focus on linear and ReLU attentions. The analysis for softmax attention follows a nearly identical proof with an additional beginning phase similar to the analysis in Section 4. Let $M$ denote the size of a dataset of $M$ i.i.d sentences $z_{H+1}$ generated from the data model in Definition 2.1. Instead of minimizing the full population loss as in existing works (Bietti et al., 2023; Huang et al., 2024a; Chen et al., 2025), which would require either knowing $\alpha$ or taking $M\to\infty$, we aim to derive a finite-sample analysis that holds for finite $M$ and unknown $\alpha$.
Before presenting our training algorithm and its finite-sample analysis, we first discuss the easier case where $\alpha$ is known and explain why it is difficult to derive the convergence rate of the population loss in (67). Observe that the function $f_{C}(\lambda,\gamma)=-C\lambda-\alpha\gamma+\ln(e^{C\lambda}+e^{C\lambda+\gamma}+N-1)$ is jointly convex and 1/2-smooth with respect to $\lambda$ and $\gamma$. Hence, at first glance, it seems that the convergence rate of $L(\lambda,\gamma)=\mathbb{E}_{C}[f_{C}(\lambda,\gamma)]$ follows from existing results in (stochastic) convex and smooth optimization (Nemirovski et al., 2009). However, as a convex function on unbounded domain with negative partial derivatives, no finite minimizer exists for $L(\lambda,\gamma)$. This implies that minimizing $L(\lambda,\gamma)$ is a multi-dimensional convex optimization problem on astral space (Dudík et al., 2022). To our knowledge, nothing is known about the convergence rate to the infimum for this problem.
@@
The empirical loss is $L_{\mathrm{emp}}(\lambda)=\frac{1}{M}\sum_{m=1}^{M}-\ln\frac{\exp(\xi_{z_{H+1}}^{(m)})}{\sum_{j\in[N+1]}\exp(\xi_{j}^{(m)})}.$ Then, we run normalized gradient descent on $L_{\mathrm{emp}}(\lambda)$ with a constant learning rate $\eta>0$. The formal procedure is given in Algorithm 1 in Appendix D.2. The following theorem shows that the population loss converges at a linear rate to the Bayes risk, similar to the noiseless setting.
@@
With probability at least $1-\delta$ over the training set of size $M$, after $t$ iterations of normalized gradient descent on $L_{\mathrm{emp}}(\lambda)$, Algorithm 1 guarantees that
$L(s=1,\lambda_{t},\hat{\gamma})\leq L_{\mathrm{Bayes}}+\frac{1}{\min(\alpha,1-\alpha)-\sqrt{\frac{\ln(2/\delta)}{2M}}}\frac{\ln(2/\delta)}{2M}+(N-1)e^{-\eta t}.$
OOD Generalization to unseen output words. Similar to Theorem 3.3 for noiseless learning, the following theorem shows that Algorithm 1 produces a trained model that generalizes to an unseen output word $y_{\mathrm{test}}\notin{\mathcal{O}}\cup{\mathcal{Q}}$ in the noisy setting. The proof can be found in Appendix D.3.
@@
$\Pr[z_{H+1}=y_{\rm test}\mid\lambda_{t},\hat{\gamma}$] $\vcentcolon=\frac{\exp(\xi_{y_{\mathrm{test}}})}{\sum_{j\in[N+1]}\exp(\xi_{j})}=1-\alpha+O\left(\sqrt{\frac{\ln(1/\delta)}{M}}+N^{2}e^{-2\eta t}\right),$
@@
The following theorem shows that (4) always hold after a sufficiently large number of iterations.
@@
Algorithm 1 guarantees that with probability $1-\delta$, the condition (4) holds after any $|t\geq\max(1,\frac{1}{\eta}\absolutevalue{\ln(1-\alpha+\sqrt{\frac{\ln(2/\delta)}{2M}})-\ln(\alpha-\sqrt{\frac{\ln(2/\delta)}{2M}})}$ training iterations.